OpenStudy (anonymous):

Going against the current, a boat takes 6 hours to make a 120-mile trip. When the boat travels with the current on the return trip, it takes 5 hours. What is the rate of the boat in still water?

OpenStudy (anonymous):

This is a system of equation problem. One sec but satelite might beat me lol

OpenStudy (anonymous):

is it 20?

OpenStudy (anonymous):

go ahead i will be quiet

OpenStudy (anonymous):

no it is not 20

OpenStudy (anonymous):

Oh man!!!

OpenStudy (anonymous):

I think I got it...44/2=22!! I did 120/5 and 120/6 added them up and divided them by 2!!

OpenStudy (anonymous):

put the rate of the boat in still water as say \(x\) and the rate of the current as \(y\)

OpenStudy (anonymous):

hold on' the average rate is seldom the average of the rates

OpenStudy (anonymous):

okay....sorry!

OpenStudy (anonymous):

we know that distance equals rate times time , i.e. \(D=R\times T\) and so \(R=\frac{D}{T}\)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

if you put \(x\) as the rate of the boat in still water and \(y\) as the rate of the current, then the combined rate with the current is \(x+y\) and against the current it is \(x-y\) you know the rate with the current is \(\frac{120}{5}=24\) so \[x+y=24\] and you know the rate against the current is \(\frac{120}{6}=20\) so \[x-y=20\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

you solve the two equations \[x+y=24\] \[x-y=20\] almost in your head but if you cannot come up with two numbers that add to 24 and whose difference is 20 then add up the two equations and get \[2x=44\] and so \[x=22\]

OpenStudy (anonymous):

therefore the boat in still water travels at \(22\) and the current is 2

OpenStudy (anonymous):

Awesome!!! Thanks!! Now would xy represent the rate of the boat going with the current?

OpenStudy (anonymous):

i do notice that you get the same answer using your method, but i assure you it is a fluke if you change to 6 hours and 7 hours it would not work this way

OpenStudy (anonymous):

It was me doing what I call monkey math!! Just trying everything and anything I could think of!! LOL

OpenStudy (anonymous):

\(x+y\) is the rate going with the current, but we already knew that. it was 24

OpenStudy (anonymous):

what you wanted was the rate of the boat in still water it is 22

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

done time for the twilight zone later

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

My take 5(x + y) = 120 // With the Current 6(x - y) = 120 / Against the Current //With the current 5(x + y) = 120 \[\frac{1}{5} * 5(x + y) = \frac{1}{5} *120 \] \[x + y = 24 \] //Against the current \[ 6(x + y) = 120 \] \[\frac{1}{6} * 6(x - y) = \frac{1}{6} *120 \] \[x - y = 20 \] //Now we have a system \[x + y = 24 \] \[x + y = 20 \] //Eliminate y and add the system \[x + y= 24 \] \[x + y = 20 \] \[2x = 44 \] \[\frac{2x} {2} = \frac{44}{2} \] \[x = 22 \] To find y, which is the current insert 22 into the equation for with the current \[22+ y = 24 \] \[ y = 24 -22 \] \[ y = 2 \] The current is 2

OpenStudy (anonymous):

Made a boo boo the system should of been \[ x+y=24 \] \[ x-y=20\]